OpenStudy (anonymous):

Hi. I'd like to know what do i do when i need to flip an equation? e.g. 3(x-2)+4x=2x^2-4x Where i need to isolate all the constants on one side?

OpenStudy (saifoo.khan):

\[3(x-2)+4x=2x^2-4x\]\[3x - 6 + 4x = 2x^2 - 4x\] \[-2x^2 + 4x + 3x - 6 + 4x = 0\] \[-2x^2 + 7x + 4x - 6=0\] \[-2x^2 + 11x - 6 = 0\] Now multiply the whole equation by -1, so that u can get the value of a as +. \[2x^2 - 11x - 6=0\]

OpenStudy (anonymous):

Than You, i felt like that was the what i needed to do but i wanted to know if there were any other procedures. Once again, thank you.